Solve the inequality and enter answer in interval notation. For example, your answer may look like (-40,inf)
19/(x+2)>=5

Answer by stanbon(75887)

Solve the inequality and enter answer in interval notation. For example, your answer may look like (-40,inf)
19/(x+2)>=5
----
1st: x cannot be -2
-----
2nd: solve the equality:
19/(x+2) = 5
19/5 = x+2
x = (19/5)-(10/5)
x = 9/5
-------
Draw a number line and plot x = -2 and x = (9/5)
---
Check values in each of the resulting intervals to
see where solutions for the inequality lie.
---
19/(x+2)>5
Let x = -10; you get 19/-8> 5; false
---
Let x = 0; you get 19/2 > 5 ; true so solutions in (-2,9/5)
---
Let x = 10; you get 19/12 > 5 ; false
----
Total solution set for the equality and the inequality;
(-2,9/5]
===============
Cheers,
Stan H.